### v1

# Resource-Efficient Federated Multimodal Learning via Layer-wise and Progressive Training

Abstract
Combining different data modalities enables deep neural networks to tackle complex tasks more effectively, making multimodal learning increasingly popular. To harness multimodal data closer to end users, it is essential to integrate multimodal learning with privacy-preserving training approaches such as federated learning (FL). However, compared to conventional unimodal learning, multimodal setting requires dedicated encoders for each modality, resulting in larger and more complex models that demand significant resources. This presents a substantial challenge for FL clients operating with limited computational resources and communication bandwidth. To address these challenges, we introduce LW-FedMML, a layer-wise federated multimodal learning approach, which decomposes the training process into multiple steps. Each step focuses on training only a portion of the model, thereby significantly reducing the memory and computational requirements. Moreover, FL clients only need to exchange the trained model portion with the central server, lowering the resulting communication cost. We conduct extensive experiments across various FL scenarios and multimodal learning setups to validate the effectiveness of our proposed method. The results demonstrate that LW-FedMML can compete with conventional end-to-end federated multimodal learning (FedMML) while significantly reducing the resource burden on FL clients. Specifically, LW-FedMML reduces memory usage by up to 2.7×2.7$\times$2.7 ×, computational operations (FLOPs) by 2.4×2.4$\times$2.4 ×, and total communication cost by 2.3×2.3$\times$2.3 ×. We also introduce a progressive training approach called Prog-FedMML. While it offers lesser resource efficiency than LW-FedMML, Prog-FedMML has the potential to surpass the performance of end-to-end FedMML, making it a viable option for scenarios with fewer resource constraints.
Index Terms:
federated learning, multimodality, layer-wise, progressive, resource-efficient.
I Introduction
Deep neural networks have become an integral part of many practical applications in our daily lives. While these models have traditionally been trained on a single modality, recent advancements in the field are driving a shift towards multimodal learning [1, 2, 3]. This transition is motivated by the realization that leveraging multiple types of data simultaneously can significantly enhance the performance of the models [4]. Just as animal intelligence relies on multiple senses to perceive their environment, combining different data modalities allows models to tackle complex tasks more effectively. Integrating diverse data sources, such as vision and audio, ensures that more information is available to the model, thereby improving its perception and decision-making capabilities [5].
Multimodal learning based on instance discrimination methods aligns representations of different modalities within the same data sample, thereby learning transferable representations [3, 6]. On the other hand, supervised multimodal learning aims to fuse information from various modalities to enhance the predictive performance of the model [7, 8, 9]. As mobile and IoT devices become increasingly powerful, they are integrated with a broader range of sensors [10]. For instance, smartphones are equipped with cameras, gyroscopes, accelerometers, and more, while smart home IoT devices can include an array of modular sensors for various functions. These edge devices capture and store valuable real-world multimodal data, which is crucial for training AI models for practical applications. However, conventional multimodal learning methods require collecting this data into a central repository for training, raising data privacy concerns.
To harness the distributed data while also considering privacy issues, numerous studies have explored decentralized training approaches like federated learning (FL) [11, 12, 13, 14]. In an FL system, the central server collects model parameters trained on the private data of edge devices (known as FL clients) rather than the data itself [11]. Such a mechanism effectively guarantees the confidentiality of data held by different parties involved in the training process. However, this also means that the FL training process relies heavily on the computation and communication capabilities of FL clients, as they are responsible for training the model [15].
While a single encoder is sufficient to process the data in unimodal tasks, multimodal tasks often require dedicated encoders tailored for each modality. Consequently, multimodal learning demands significant computational and memory resources to accommodate these modality-specific encoders. This presents a drawback for FL clients, which are often edge devices with limited resources, making it difficult to train multimodal models [16]. The problem becomes even more challenging when dealing with high-dimensional data, such as video or audio.
The need for modality-specific encoders makes multimodal models larger and more complex than their unimodal counterparts, leading to increased transmission costs. This creates communication bottlenecks in the FL process, as FL clients may struggle to exchange these large multimodal models with the central server during training. Due to diverse hardware capabilities and resource constraints, many FL clients may not meet the resource requirements for conventional end-to-end multimodal training. As a result, their valuable data is excluded from the training process. Therefore, exploring resource-efficient approaches is crucial for maximizing the effectiveness of multimodal learning within the FL context.
Figure 1: Overview of the federated learning process. (i) The global model $M$ is distributed to clients. (ii) Each client performs training on the local dataset $D^{n}$. (iii) The local model $M^{n}$ is sent back to the server. (iv) The server performs aggregation.
During training, reducing the batch size may appear to be a straightforward method to lower the computational demands on resource-constrained devices. However, for instance discrimination-based multimodal learning approaches (e.g., CLIP [3]), a small batch size may compromise the effectiveness of the contrastive loss, degrading the quality of learned representations [17, 18]. These methods benefit from a larger batch size to increase the number of negative samples, which is essential for the contrastive loss to perform well. Furthermore, simply reducing the batch size only addresses the computational requirements but fails to alleviate the communication burden on FL clients, as they still need to exchange the same model size with the server.
To address the resource constraints in an FL environment for multimodal learning, our work introduces the layer-wise training approach. This approach decomposes the training process into multiple stages, with each stage dedicated to training a single layer or a portion of the model. By training different layers in separate stages, we can effectively reduce the computational and memory requirements on FL clients. Furthermore, at each stage, only the layer active for training needs to be exchanged with the server, thereby reducing the communication burden on FL clients.
Our contributions are summarized as follows: (i) We introduce LW-FedMML, a layer-wise federated multimodal learning strategy designed to address the resource limitations of FL clients engaged in multimodal tasks. LW-FedMML can match the performance of conventional end-to-end training while significantly reducing resource requirements. (ii) We also explore a progressive training approach, referred to as Prog-FedMML. Although Prog-FedMML does not offer the same level of resource efficiency as LW-FedMML, it has the potential to outperform conventional end-to-end training while maintaining similar memory usage and lowering computational and communication costs. (iii) To validate the effectiveness of our proposed methods, we conducted extensive experiments across various FL scenarios and multimodal learning setups.
(a) Instance discrimination-based multimodal learning.
(b) Supervised multimodal learning.
Figure 2: Common multimodal learning approaches with modalities $a$ and $b$. $M$ denotes the model, and $x$ denotes the input sample. (a) The instance discrimination-based approach focuses on aligning the representations of diverse modalities within a given data sample. Negative samples are omitted in the figure for clarity. (b) The supervised approach aims to fuse information from various modalities, enabling the model to make more informed predictions. Here, $H_{\text{sup}}$ can be any task-specific prediction head.
Figure 3: Overview of LW-FedMML across different stages. In stage $s$, the active layers $L_{a}^{s}$ and $L_{b}^{s}$ are depicted in green within encoders $F_{a}$ and $F_{b}$, where $a$ and $b$ represent different modalities. Prior layers within the encoders are frozen, as indicated in gray.
Figure 4: Overview of LW-FedMML for supervised setting at stage $s$.
II Preliminary
II-A Federated Learning
A federated learning (FL) process typically involves a central server and a set of FL client devices [11, 19, 20]. These devices are often found close to end-users, capturing and holding valuable data for model training. Meanwhile, the central server oversees and monitors the overall FL procedure. Given $N$ clients, the goal of FL is to train a global model $M$ by solving the following optimization problem:
$M^{*}=\operatorname*{arg\,min}_{M}\sum_{n=1}^{N}w^{n}\mathcal{L}(M,D^{n}),$ (1)
where $\mathcal{L}$ denotes the loss function and $D^{n}$ represents the local dataset of the $n$-th client. Here, the weight $w^{n}$ assigned to client $n$ is typically calculated as:
$w^{n}=\frac{|D^{n}|}{|D|}\ \text{,}\ \text{where}\ {D=\bigcup_{n=1}^{N}D^{n}},$ (2)
with $|D^{n}|$ and $|D|$ representing the number of data samples in $D^{n}$ and $D$, respectively. As illustrated in Fig. 1, four key steps comprise a standard FL communication round. (i) The central server distributes the global model $M$ to each participating client $n\in[1,N]$. (ii) Each client $n$ trains the received model on its local dataset $D^{n}$ to produce the local model $M^{n}$. (iii) The trained local models $[M^{1},\dots,M^{N}]$ are sent back to the server. (iv) The server aggregates the received local models and updates the global model. The FedAvg algorithm is commonly used for the aggregation process, which can be described as $M\leftarrow\sum^{N}_{n=1}w^{n}M^{n}$. These four steps are iteratively executed until the model converges or a predefined number of FL communication rounds $R$ is reached.
II-B Multimodal Learning
In many complex tasks, such as healthcare, autonomous driving, and weather forecasting, a single modality may not provide sufficient context or detail for accurate decision-making [21, 22, 23]. Multimodal learning is essential in these scenarios, as it leverages the complementary information of different data modalities, providing a more comprehensive and detailed understanding of the task at hand. Many studies on multimodal learning center around aligning the representations of diverse modalities within a given data sample, as illustrated in Fig. 2a. These approaches often employ contrastive loss, as seen in CLIP [3], to draw representations of different modalities within the same sample closer together, while simultaneously pushing those from different samples further apart. This allows encoders of different modalities to learn transferable representations applicable to a wide range of downstream tasks. These pretrained encoders can then be deployed either individually or in pairs, depending on whether the downstream tasks involve unimodality or multimodality processing. On the other hand, supervised multimodal learning, depicted in Fig. 2b, seeks to fuse information from various modalities, enabling the model to make more accurate and informed predictions [7, 8, 9]. A straightforward way to fuse representations of different modalities is to concatenate them, thereby creating a unified feature representation that contains the essence of each modality. However, supervised approaches rely on labeled data for training, which can be costly and labor-intensive to obtain, often serving as a limiting factor in many cases.
III Related Work
Many studies have shown increasing interest in federated multimodal learning due to its potential to be more powerful than conventional unimodal approaches [16]. In [24], the authors propose a co-attention mechanism to fuse information from different modalities within clients. Hierarchical gradient blending is introduced in [25] to address the challenge of varying generalization rates among modality sub-networks and local models. The authors in [26] aim to learn fine-grained image representations from vision and language modalities in the FL setting. In the medical field, [27] demonstrates the application of federated multimodal learning to detect melanoma from skin lesion images and clinical data. Similarly, [28] utilizes CT and MRI images for liver segmentation. These studies highlight the versatility and potential of federated multimodal learning to leverage diverse data sources in various domains.
The layer-wise training approach was originally employed to train deep belief networks and restricted Boltzmann machines [29, 30], to avoid training issues such as vanishing gradients. This approach can significantly reduce resource requirements in the context of federated multimodal learning, where the integration of multiple modalities imposes even greater computational and communication demands compared to unimodal FL. The authors in [31] propose a progressive training approach that incrementally trains the model by increasing its depth. Similarly, [32] and [33] explore layer-wise training approaches for federated self-supervised learning. In contrast, these aforementioned studies primarily consider unimodal settings. Our work extends layer-wise and progressive training approaches to the multimodal context in FL, addressing the unique challenges and opportunities presented by multimodal data in FL environments.
Algorithm 1 LW-FedMML (Server-side)
Input: encoder $F_{a}^{0}$ and projection head $H_{a}$ for modality $a$, encoder $F_{b}^{0}$ and projection head $H_{b}$ for modality $b$, number of stages $S$, number of rounds $R^{s}$ where $s\in[1,S]$
Output: $F_{a}^{S}$, $F_{b}^{S}$, $H_{a}$, $H_{b}$
1:Server executes:
2:for stage $s=1,2,\dots,S$ do
3:     Initialize new layers: $L_{a}^{s}$ and $L_{b}^{s}$
4:     for round $r=1,2,\dots,R^{s}$ do
5:         for client $n=1,2,\dots,N$ in parallel do
6:              $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}\leftarrow\text{Train}(L_{a}^{s},L_{b}^{s},H_{a},H_{b})$
7:              $w^{n}=\frac{|D^{n}|}{|D|}$, where $D=\bigcup_{n=1}^{N}D^{n}$
8:         end for
9:         $L_{a}^{s}\leftarrow\sum_{n=1}^{N}w^{n}L_{a}^{s,n},\quad H_{a}\leftarrow\sum_{n=1}^{N}w^{n}H_{a}^{n}$
10:         $L_{b}^{s}\leftarrow\sum_{n=1}^{N}w^{n}L_{b}^{s,n},\quad H_{b}\leftarrow\sum_{n=1}^{N}w^{n}H_{b}^{n}$
11:     end for
12:     Distribute $L_{a}^{s}$ and $L_{b}^{s}$ to clients
13:     $F_{a}^{s}\leftarrow$ Attach $L_{a}^{s}$ to $F_{a}^{(s-1)}$
14:     $F_{b}^{s}\leftarrow$ Attach $L_{b}^{s}$ to $F_{b}^{(s-1)}$
15:end for
16:return $F_{a}^{S}$, $F_{b}^{S}$, $H_{a}$, $H_{b}$
IV Proposed Method
IV-A LW-FedMML
The main idea of layer-wise training is to break down the training process of a model into multiple stages, denoted as $s\in[1,S]$. Initially, a given encoder $F$ starts off without any layers, essentially as an empty framework. At the beginning of each stage $s$, a new layer $L^{s}$ is sequentially attached to the encoder $F$, thereby increasing its depth. During stage $s$, only the corresponding layer $L^{s}$ is trained, while the other layers $[L^{1},\dots,L^{(s-1)}]$ from previous stages are frozen without receiving updates. Each stage $s$ lasts for a specified number of epochs, or communication rounds in the context of FL.
Fig. 3 illustrates the training process for LW-FedMML. Specifically, given different modalities $a$ and $b$, layers $L^{s}_{a}$ and $L^{s}_{b}$ are attached to encoders $F_{a}$ and $F_{b}$ at the beginning of each stage $s$. For LW-FedMML without labels, we follow the CLIP [3] approach, which computes the InfoNCE [34] loss for each modality. This approach allows for effective alignment of representations from different data modalities with simplicity. Give a batch of samples $x_{a}$ and $x_{b}$, the projections are extracted as $z_{a}=H_{a}(F_{a}(x_{a}))$ and $z_{b}=H_{b}(F_{b}(x_{b}))$, where $H_{a}$ and $H_{b}$ denote modality-specific projection heads. When both $z_{a}$ and $z_{b}$ are normalized, the loss for modality $a$ can be calculated as follows:
$\ell(z_{a},z_{b})=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(z_{a}^{i}\cdot z_{b}^{i}/\tau)}{\sum_{j=1}^{B}\exp(z_{a}^{i}\cdot z_{b}^{j}/\tau)},$ (3)
where $B$ is the batch size and $\tau$ represents the temperature parameter. This loss is designed to promote the cosine similarity between representations of different modalities originating from the same sample. Simultaneously, it pushes the representations from different samples away from each other. Following the same principles, the loss for modality $b$ can be calculated as $\ell(z_{b},z_{a})$. The final training loss can be defined as:
$\ell_{clip}(z_{a},z_{b})=\frac{1}{2}(\ell(z_{a},z_{b})+\ell(z_{b},z_{a})).$ (4)
In supervised LW-FedMML, the concatenated representations of $z_{a}$ and $z_{b}$ are passed through a prediction head $H_{\text{sup}}$ for making predictions, as shown in Fig 4. The cross-entropy loss is calculated between the prediction and the ground truth label for training.
Algorithm 2 LW-FedMML (Client-side)
Input: local dataset $D^{n}$, layers $[L_{a}^{1},\dots,L_{a}^{(s-1)}]$ for modality $a$, layers $[L_{b}^{1},\dots,L_{b}^{(s-1)}]$ for modality $b$, number of local epochs $E$
Output: $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}$
1:Client executes: $\text{Train}(L_{a}^{s},L_{b}^{s},H_{a},H_{b})$:
2:$L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}\leftarrow$ Download $L_{a}^{s},L_{b}^{s},H_{a},H_{b}$
3:$F_{a}^{s,n}\leftarrow$ Freeze $[L_{a}^{1},\dots,L_{a}^{(s-1)}]$ and attach $L_{a}^{s,n}$
4:$F_{b}^{s,n}\leftarrow$ Freeze $[L_{b}^{1},\dots,L_{b}^{(s-1)}]$ and attach $L_{b}^{s,n}$
5:for epoch $e=1,2,\dots,E$ do
6:     for each batch $(x_{a},x_{b})\in D^{n}$ do
7:         $z_{a}\leftarrow H_{a}^{n}(F_{a}^{s,n}(x_{a}))$
8:         $z_{b}\leftarrow H_{b}^{n}(F_{b}^{s,n}(x_{b}))$
9:         $\mathcal{L}\leftarrow\ell_{\text{clip}}(z_{a},z_{b})$
10:         $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}\leftarrow$ Update with $\nabla\mathcal{L}$
11:     end for
12:end for
13:Upload $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}$
TABLE I: Comparison of layer management strategies across different training approaches.
At Stage $\mathbf{s\in[1,S]}$
Training
Number of
Stages
Starting
Layers
Existing
Layers
Added
Layer
Frozen
Layers
Active
Layers
FedMML end-to-end $L^{1}$ to $L^{S}$ $L^{1}$ to $L^{S}$ $L^{1}$ to $L^{S}$
Prog-FedMML progressive $S$ $L^{1}$ to $L^{s}$ $L^{s}$ $L^{1}$ to $L^{s}$
LW-FedMML layer-wise $S$ $L^{1}$ to $L^{s}$ $L^{s}$ $L^{1}$ to $L^{(s-1)}$ $L^{s}$
TABLE II: Distribution of FL communication rounds for each stage of layer-wise training on the COCO dataset.
Rounds $\boldsymbol{R_{s}}$ for Stage $\boldsymbol{s}$ Total Number of Rounds Text-to-Image Image-to-Text
1 2 3 4 5 6 R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
LW-FedMML 15 15 15 15 15 15 90 21.26 48.54 62.21 89.30 30.48 59.32 71.58 92.50
LW-FedMML 0 0 0 10 30 50 90 25.22 53.82 66.92 91.30 35.52 65.78 76.88 94.20
TABLE III: Learning rate for encoders of different modalities.
Dataset Modality Stage $\boldsymbol{s}$
1 2 3 4 5 6
COCO Image 1e-7 1e-7 1e-7 5e-6 3e-5 5e-5
Text 1e-7 1e-7 1e-7 1e-5 3e-5 5e-5
ADVANCE Image 1e-7 1e-7 1e-7 5e-7 1e-6 5e-6
Audio 1e-5 1e-5 1e-5 1e-5 1e-5 3e-5
FL clients perform the aforementioned steps as the local training process. The detailed procedure of LW-FedMML is shown in Algorithms 1 and 2. Algorithm 1 describes the server-side execution, responsible for determining the stage $s$ and initializing new layers $L_{a}^{s}$ and $L_{b}^{s}$ for each modality, as well as completing the aggregation process. On the other hand, the clients execute Algorithm 2, where they train the active layers $L_{a}^{s}$ and $L_{b}^{s}$ along with the projection heads $H_{a}$ and $H_{b}$. For the supervised setting, clients also train the prediction head $H_{\text{sup}}$ and exchange it with the server. By training only the active layers in each stage, memory and computational resource requirements for clients are significantly reduced. This enables resource-limited client devices to actively participate in the FL process, allowing to take advantage of their private data for model training. Furthermore, only the active layers are required to exchange between the clients and the server. Consequently, both upload and download costs for model exchange are reduced, minimizing FL communication bottlenecks.
IV-B Prog-FedMML
Inspired by [31], we also explore the progressive training approach for federated multimodal learning, denoted as Prog-FedMML. Progressive training, as demonstrated in [31] and [32], has the potential to outperform end-to-end training. Similar to the layer-wise approach, progressive training begins with an initial encoder $F$ without any layers. A new layer $L^{s}$ is sequentially added at each stage $s$. However, unlike the layer-wise approach, all existing layers in $F$ are trained during each stage $s$. Specifically for multimodal learning, Prog-FedMML trains layers $[L_{a}^{1},\dots,L_{a}^{s}]$, and $[L_{b}^{1},\dots,L_{b}^{s}]$ within encoders $F_{a}$ and $F_{b}$ during stage $s$. As a result, Prog-FedMML is more resource-intensive than LW-FedMML, as it requires training all layers added in prior stages, rather than freezing them. The computational and communication resource requirements of Prog-FedMML for an FL training round gradually increase and eventually match those of end-to-end training by the final stage (i.e., $s=S$). This similarity in resource demands can lead to the same drawbacks as end-to-end training, particularly in excluding clients with limited resources from the FL process, thereby losing valuable data for model training. We compare the different characteristics of conventional end-to-end, progressive, and layer-wise training approaches for an encoder $F$ in Table I, highlighting how the layers are managed and utilized at each stage.
Figure 5: For training on the COCO dataset, we attach two transformer blocks to the ViT-Tiny encoder and one transformer block to the DistilBERT encoder at each stage. Similarly, for the ADVANCE dataset, we attach two transformer blocks to the ViT-Tiny encoder and one transformer block to the DistilAST encoder at each stage.
TABLE IV: Comparison between FedMML (end-to-end), Prog-FedMML (progressive), and LW-FedMML (layer-wise) for text-to-image and image-to-text retrieval tasks on the COCO dataset. We also compare the relative resource requirements of each approach for a client.
Resource Requirements Text-to-Image Image-to-Text
Memory FLOPs Comm. R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 26.31 55.06 67.91 91.58 37.48 66.00 78.12 94.52
Prog-FedMML 0.95$\times$ 0.59$\times$ 0.82$\times$ 27.42 56.14 69.38 92.77 39.02 68.32 79.10 95.54
LW-FedMML 0.37$\times$ 0.42$\times$ 0.43$\times$ 25.22 53.82 66.92 91.30 35.52 65.78 76.88 94.20
(a) Memory.
(b) FLOPs.
(c) Communication.
Figure 6: Computational and communication resources required for a client on the COCO dataset. (a) Memory requirement. (b) FLOPs consumption. (c) Communication cost.
V Experiment
V-A Experimental Setup
Dataset. We conduct our experiments for multimodal learning in two settings: an instance discrimination-based setting and a supervised setting with labeled data. We use the COCO [35] dataset for the instance discrimination-based setting. This dataset contains 118,287 images for training and 5,000 images for testing, with each image paired with at least five captions. We evaluate the performance on text-to-image and image-to-text retrieval tasks. For the supervised approach, we use the ADVANCE [9] dataset for audiovisual aerial scene recognition. This dataset consists of 5,075 aerial images accompanied by their respective audio recordings, categorized into 13 different scene classes. The underlying premise is that geographic locations are associated with certain sound events, and utilizing both aerial images and sound can lead to more accurate predictions. We allocate 20% of the data for testing.
Model. We choose ViT-Tiny [36] for the image modality, DistilBERT [37] for the text modality, and a distilled version of the audio spectrogram transformer [38], which we denote as DistilAST, for the audio modality. We utilize pre-trained weights from the HuggingFace Transformers library [39] for these models. It is important to highlight that these pre-trained weights are derived from training separately on unimodal data, and not on multimodal data or tasks. ViT-Tiny is paired with DistilBERT for the COCO dataset, and with DistilAST for the ADVANCE dataset. The projection head $H$ is implemented as a 3-layer MLP for all modalities, with the dimension of hidden layers set to 4096 and the output dimension set to 256. For the supervised setting, the prediction head $H_{\text{sup}}$ is a 2-layer MLP with hidden layer dimensions of 256.
Training settings. For layer-wise training, we break down the training process into six stages. In this context, the term “layer” refers to a transformer block or a model portion that comprises multiple layers. The ViT-Tiny architecture consists of 12 transformer blocks, while both DistilBERT and DistilAST contain six transformer blocks. For the COCO dataset, we attach two transformer blocks to the ViT-Tiny encoder and one transformer block to the DistilBERT encoder at each training stage, as illustrated in Fig. 5. Similarly, for the ADVANCE dataset, we attach two transformer blocks to the ViT-Tiny encoder and one transformer block to the DistilAST encoder at each stage, following the same training procedure. We set the number of FL clients to 10, with each client performing local training for three epochs ($E=3$). The total number of FL communication rounds $R$ is set to 90.
We uniformly split the data among the clients for the COCO dataset. We set the batch size to 256 and use $\tau=0.05$ in Eq. 3. For LW-FedMML and Prog-FedMML, we use the learning rates shown in Table III for the image and text encoders at each stage. The learning rate for the projection heads of both modalities is set to 3e-4. For the conventional end-to-end FedMML, the learning rates for the image and text encoders are set to 1e-7 and 1e-5, respectively, while the projection heads have a learning rate of 1e-4. Unless otherwise stated, we use the COCO dataset for our experiments.
[…]
TABLE V: Performance comparison on the COCO dataset, assuming that only a portion of clients can meet the resource requirements for end-to-end training (FedMML) and progressive training (Prog-FedMML) approaches.
Resource Requirements Eligible Clients Text-to-Image Image-to-Text
Memory
FLOPs
(×$\bm{(\times 10^{12})}$
Comm.
R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
FedMML 19913 2.44 76613 5/10 (50%) 22.48 49.03 62.47 88.50 30.02 59.44 72.30 92.10
3/10 (30%) 18.45 43.98 57.18 84.66 25.30 52.48 66.34 90.14
Prog-FedMML 18969 1.45 62918 5/10 (50%) 23.89 52.22 65.27 90.94 34.38 63.08 74.78 93.70
3/10 (30%) 19.48 46.08 59.82 87.24 26.76 56.82 69.02 91.44
LW-FedMML 7371 1.03 32794 10/10 (100%) 25.22 53.82 66.92 91.30 35.52 65.78 76.88 94.20
(a) Text-to-Image.
(b) Image-to-Text.
(c) Memory usage.
Figure 7: Comparison with different batch sizes on the COCO dataset.
V-B Instance Discrimination-based Setting
Given a limited number of FL rounds (i.e., $R=90$), we first determine how to allocate them to different stages for layer-wise training. As shown in Table II, we experiment with two strategies: uniformly distributing the FL rounds across all stages and allocating more rounds to the later stages. Prior work [32] indicates that allocating the FL rounds uniformly or more to earlier stages is effective in the absence of pre-trained weights. However, our findings in Table II demonstrate that allocating more rounds to the later stages achieves superior performance compared to uniform distribution when using pretrained weights. Therefore, we adopt this allocation strategy in our subsequent experiments. For Prog-FedMML, we use uniform allocation unless otherwise stated.
In Table IV, we compare the performance of different approaches and the relative resource requirements for a single client. We measure the maximum memory required during training, while FLOPs and communication costs are measured in total across all local training epochs and FL rounds. For FLOPs calculation, we follow the same strategy as in [32], accounting for only one input sample rather than the entire batch. As the number of samples increases, the FLOPs would increase drastically. Table IV demonstrates that LW-FedMML achieves comparable performance to end-to-end FedMML, with a minor drop of $\approx 1\%$ in average recall for text-to-image and image-to-text retrieval tasks, while significantly reducing resource requirements. Specifically, LW-FedMML shows a 2.7×2.7$\times$2.7 × reduction in memory usage, a 2.4×2.4$\times$2.4 × decrease in FLOPs requirements, and 2.3×2.3$\times$2.3 × lower communication costs. Prog-FedMML achieves the best performance, surpassing end-to-end FedMML, which aligns with findings from prior studies [31, 32]. However, the resource requirements of Prog-FedMML are significantly higher than those of LW-FedMML, closely matching those of end-to-end FedMML in terms of memory and communication costs. This could present similar challenges for Prog-FedMML as those faced by end-to-end FedMML, where many clients may fail to meet the resource requirements for participation, thereby preventing their data from being used in model training. Fig. 6 plots the memory usage, FLOPs, and communication costs during the training process for a client. Across all metrics, LW-FedMML consistently shows significantly lower resource requirements compared to FedMML. In contrast, resource demands for Prog-FedMML gradually increase throughout the training, eventually matching those of FedMML.
(a) FedMML
(b) Prog-FedMML
Figure 8: Relative gradient values of linear layers corresponding to the attention output in DistilBERT. (a) In FedMML, the gradient value diminishes as it backpropagates to earlier layers. (b) Conversely, Prog-FedMML maintains improved gradient flow, allowing for more effective updates.
TABLE VI: Comparison between FedMML (end-to-end), Prog-FedMML (progressive), and LW-FedMML (layer-wise) for audiovisual aerial scene recognition on the ADVANCE dataset. We also compare the relative resource requirements of each approach for a client.
Resource Requirements
Memory FLOPS Comm. Precision Recall F1
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 99.82 94.49 96.93
Prog-FedMML 1.00$\times$ 0.91$\times$ 0.95$\times$ 100.00 95.76 97.75
LW-FedMML 0.51$\times$ 0.41$\times$ 0.54$\times$ 99.90 95.17 97.36
(a) Memory.
(b) FLOPs.
(c) Communication.
Figure 9: Computational and communication resources required for a client on the ADVANCE dataset. (a) Memory requirement. (b) FLOPs consumption. (c) Communication cost.
V-C Tracking Gradients
We investigate the potential cause behind why progressive training outperforms conventional end-to-end training. Using the text encoder (i.e., DistilBERT) with six transformer blocks, we examine the gradient values of linear layers corresponding to the attention output in each transformer block. For Prog-FedMML, we measure the gradient value for each layer at its corresponding stage. The gradient values relative to the layer in the last transformer block are illustrated in Fig. 8. For FedMML, the gradient value diminishes as it backpropagates to earlier layers. This diminishing pattern is not observed in Prog-FedMML. Although not yet thoroughly investigated, increasing the model depth step-by-step may enable earlier layers to receive more effective gradient updates by mitigating small updates caused by vanishing gradients or large divergences due to exploding gradients. We believe this is a key factor contributing to the superior performance of progressive training compared to conventional end-to-end training.
V-D Eligible Clients
Due to the resource-constrained nature of FL clients, not all clients may be eligible to perform end-to-end or progressive training, which have high resource demands. This means that only a subset out of all available clients can contribute to FedMML and Prog-FedMML. Consequently, the valuable data on these low-resource clients cannot be used, adversely affecting overall performance. Table V compares this scenario, showing that as the number of clients eligible for training decreases, the performance of FedMML and Prog-FedMML also declines. In contrast, LW-FedMML, with its lower resource requirements, has the potential to allow these clients to participate in the training. This inclusion enables the use of data from low-resource devices, thereby improving the overall effectiveness of the training process.
(a) $\beta=50$
(b) $\beta=5$
(c) $\beta=0.5$
(d) $\beta=0.05$
Figure 10: Data distribution among clients with different $\beta$ values for the ADVANCE dataset. A darker box represents a higher number of samples for a data class within a client.
TABLE VII: Performance comparison across varying levels of data heterogeneity on the ADVANCE dataset.
$\beta=50$ $\beta=5$ $\beta=0.5$ $\beta=0.05$
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
FedMML 99.92 95.07 97.32 99.92 95.07 97.32 99.82 94.49 96.93 96.70 74.17 81.76
Prog-FedMML 100.00 96.49 98.13 100.00 96.43 98.11 100.00 95.76 97.75 95.31 76.17 82.80
LW-FedMML 100.00 95.42 97.54 99.92 95.41 97.49 99.90 95.17 97.36 96.60 74.51 81.36
TABLE VIII: Performance comparison on the ADVANCE dataset with varying numbers of participating clients $N_{\text{p}}$ at each FL round. The total number of clients $N$ is set to 40, and the total number of FL communication rounds $R$ is set to 120.
$N_{\text{p}}/N$ $30/40$ $(75\%)$ $20/40$ $(50\%)$ $10/40$ $(25\%)$
Precision Recall F1 Precision Recall F1 Precision Recall F1
FedMML 99.87 93.27 96.29 99.87 93.47 96.39 99.87 93.71 96.50
Prog-FedMML 99.95 92.93 96.08 99.90 92.84 95.98 99.95 93.27 96.22
LW-FedMML 99.90 90.44 94.59 99.90 90.54 94.62 99.90 90.54 94.62
V-E Batch Size
Even though a smaller batch size can reduce memory usage, instance discrimination-based approaches benefit from larger batch sizes due to the increased number of samples. We compare the performance of different training approaches with various batch sizes in Figs. 7a and 7b, while Fig. 7c illustrates the maximum memory usage for each setting. The results indicate that smaller batch sizes lead to reduced performance across all methods. Fig. 7c shows a sharp increase in memory requirements for end-to-end FedMML and Prog-FedMML even as the batch size grows from 32 to 256. In contrast, the memory requirements for LW-FedMML remain relatively low across the same batch sizes. Fig. 7 also demonstrates that LW-FedMML can achieve the same performance as end-to-end FedMML when using a larger batch size (i.e., 1024) while still using less memory than FedMML with a batch size of 256. Moreover, increasing the batch size in LW-FedMML only impacts the memory usage, allowing it to maintain its advantage of lower communication costs.
V-F Supervised Setting
For the ADVANCE dataset, we allocate FL rounds to later stages for Prog-FedMML, as we do for LW-FedMML in Table II. Table VI compares the performance of different approaches and relative resource requirements for a single client. The measurement of resource requirements follows the same approach as in the instance discrimination-based setting. The table shows that Prog-FedMML achieves the best performance, while LW-FedMML maintains competitive performance across various metrics compared to FedMML in a supervised multimodal learning setting. In Fig. 9, we can observe that LW-FedMML significantly lowers the resource requirements compared to FedMML, while Prog-FedMML exhibits similar resource demands as FedMML.
V-G Data Heterogeneity
Data heterogeneity is one of the main challenges in FL [20]. Variations in device usage patterns, geographical locations, and other factors influence the data distribution among clients. To control the strength of data heterogeneity within the clients’ local datasets, we adjust the $\beta$ parameter in the Dirichlet distribution. A smaller $\beta$ value corresponds to a higher level of data heterogeneity, and vice versa. Fig. 10 illustrates the data distribution among clients with different $\beta$ values for the ADVANCE dataset, and Table VII shows their results. These results indicate that the performance of all approaches declines as the level of data heterogeneity increases. Nonetheless, in each setting, Prog-FedMML achieves the best performance for most metrics, while LW-FedMML consistently maintains good performance comparable to the end-to-end FedMML.
V-H Client Participation Ratio
In an FL environment with $N$ total clients, it is often impractical for all $N$ clients to simultaneously participate in every FL training round. Therefore, at each FL round, we randomly select a subset of clients to participate in the training. For this experiment, we fixed the total number of clients to $N=40$ and set the number of participating clients $N_{\text{p}}$ to be 10, 20, and 30, while using the ADVANCE dataset. The total number of FL communication rounds $R$ is set to 120. For both LW-FedMML and Prog-FedMML, we allocate the FL rounds for each stage $R_{s}$ as $[0,0,0,20,30,70]$. The results in Table VIII demonstrate that LW-FedMML can sustain reasonable performance even when the client participation ratio is low. Prog-FedMML achieves slightly higher precision but shows a minor drop in recall and F1 metrics compared to FedMML.
VI Conclusion
Resource constraints are a significant challenge in the FL process, mainly because FL clients are often edge devices with limited capabilities. This challenge is exacerbated in FL scenarios involving multimodal data due to the complexity of the models required for processing such data. Many clients cannot meet the resource requirements for conventional end-to-end training approaches in these scenarios, resulting in their valuable data being excluded from the training process. To address this issue, we introduce LW-FedMML, a layer-wise training approach designed to significantly reduce the resource burden on clients. Our experiments demonstrate that LW-FedMML can compete with conventional end-to-end FedMML while achieving up to 2.7×2.7$\times$2.7 × lower memory usage, 2.4×2.4$\times$2.4 × FLOPs requirements, and 2.3×2.3$\times$2.3 × cheaper communication costs on the COCO dataset. These findings highlight the potential of LW-FedMML to enable efficient and scalable multimodal learning in resource-constrained FL environments. Additionally, we explore Prog-FedMML, a progressive training approach for federated multimodal learning. While Prog-FedMML has higher resource requirements than LW-FedMML, it can outperform end-to-end FedMML in most cases.